OpenStudy (hy123):

Solve the following differential equation: y' = (y^2-1)/(2t)

ganeshie8 (ganeshie8):

separate variables

OpenStudy (hy123):

Yes.

OpenStudy (hy123):

I know how to do it until i got to one point and blanked out.

ganeshie8 (ganeshie8):

dy/(y^2-1) = dt/(2t)

OpenStudy (hy123):

Yeah. I then multiplied both sides by 2

ganeshie8 (ganeshie8):

for left side, consider trig substitution \(y = sin \theta\)

OpenStudy (hy123):

But should i multiply both sides by two or it doesn't matter?

ganeshie8 (ganeshie8):

it doesnt matter

OpenStudy (hy123):

mmk.

ganeshie8 (ganeshie8):

just take the integral both sides by any means thats possible to u..

OpenStudy (hy123):

I'll take it using your way.

OpenStudy (hy123):

Can't i just use ln |y^2-1| as the integrate of the left side?

OpenStudy (hy123):

and the right side be ln |2t|

ganeshie8 (ganeshie8):

lets see

ganeshie8 (ganeshie8):

\(\large \frac{dy}{y^2-1} = \frac{ dt}{2t} \) integrate both sides \(\large \int \frac{dy}{y^2-1} = \int \frac{ dt}{2t} \) \(\large \int \frac{dy}{y^2-1} = \frac{1}{2}\int \frac{ dt}{t} \) \(\large \int \frac{dy}{y^2-1} = \frac{1}{2}\ln |t | + c \)

ganeshie8 (ganeshie8):

right side, u can pull out constant out of the integral... u still need to deal wid integrating left side

OpenStudy (hy123):

Right. But wouldnt left side be basically: